# On the asymptotics of orthogonal polynomials on multiple intervals with non-analytic weights

### v2

@@ 2.1. The first $\overline{\partial}$ deformation of the FIK RH problem
We will use this factorization locally near each interval $[a_{j},b_{j}]$. Define
$\mathbf{Z}_{n}(x,y)=\mathbf{Y}_{n}(z;\mu)\begin{cases}\begin{bmatrix}1&0\\ -f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,+},\vspace{3pt}\\ \vspace{3pt}\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,-},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
@@
¯$\overline{\partial}\mathbf{Z}_{n}(x,y)=\begin{cases}\mathbf{Y}_{n}(z;\mu)\begin{bmatrix}0&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,+},\vspace{3pt}\\ \vspace{3pt}\mathbf{Y}_{n}(z;\mu)\begin{bmatrix}0&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,-},\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}).\end{cases}$
@@
¯$\overline{\partial}\mathbf{Z}_{n}(x,y)=\begin{cases}\mathbf{Z}_{n}(z;\mu)\begin{bmatrix}0&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,+},\vspace{3pt}\\ \vspace{3pt}\mathbf{Z}_{n}(z;\mu)\begin{bmatrix}0&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,-},\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}).\end{cases}$
@@
$\mathbf{Z}_{n}^{+}(x,y)=\mathbf{Z}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j},\vspace{3pt}\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in[a_{j},b_{j}].\end{cases}$
@@
We find that $\mathbf{S}_{n}(x,y)$ solves the normalized hybrid $\overline{\partial}$ problem
$\mathbf{S}_{n}^{+}(x,y)$ $=\mathbf{S}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&1\end{bmatrix}&z\in\Sigma_{j},\vspace{3pt}\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in(a_{j},b_{j}),\vspace{3pt}\\ \,\mathrm{e}^{-n\Delta_{j}{\boldsymbol{\sigma}_{3}}}&z\in(b_{j},a_{j+1}),\end{cases}$
¯$\overline{\partial}\mathbf{S}_{n}(x,y)$ $=\mathbf{S}_{n}(z)\begin{cases}\begin{bmatrix}0&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&0\end{bmatrix}&z\in\Omega_{j,+},\vspace{3pt}\\ \vspace{3pt}\begin{bmatrix}0&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&0\end{bmatrix}&z\in\Omega_{j,-},\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}),\end{cases}$
@@
We recall here that $\mathfrak{g}_{+}(x)+\mathfrak{g}_{-}(x)=0$ for $x\in[a_{j},b_{j}]$ and $\mathfrak{g}_{+}(x)-\mathfrak{g}_{-}(x)={\Delta}_{j}$ for $x\in[b_{j},a_{j+1}]$ where $\Delta_{j}$ is a constant.
@@ 2.2. The second $\overline{\partial}$ deformation of the FIK RH problem
$d_{j,-1}(x;c)$ $=\left[\sqrt{\frac{f_{j}(x;c)}{f_{j}(x)}}-1\right]b_{\tau}(x;c),\quad c=a_{j},b_{j},$
$d_{j,1}(x;c)$ $=\left[\sqrt{\frac{f_{j}(x)}{f_{j}(x;c)}}-1\right]b_{\tau}(x;c),\quad c=a_{j},b_{j},$
and set $d_{j,\pm}^{\sharp}(x,y;c)$ to be the extension, as in Appendix B, with $\tau=\delta<\epsilon/4$ and $\kappa=0$$d_{j,\pm}^{\sharp}(x,y;c)=d_{j,\pm}(x;c)$ for $|y|<\epsilon/4$. We write
@@
for $\xi$ between $f_{j}(x;c)$ and $f_{j}(x)$. Since both of these functions are bounded below on $[0,2\delta]$ we find
$\|d_{j,\pm 1}(\cdot;a_{j})\|_{C^{0}([a_{j},b_{j}])}$ $\leq C\|f_{j}(\cdot;a_{j})-f_{j}(\cdot)\|_{C^{0}([a_{j},a_{j}+2\delta])},$
$\|d_{j,\pm 1}(\cdot;b_{j})\|_{C^{0}([a_{j},b_{j}])}$ $\leq C\|f_{j}(\cdot;a_{j})-f_{j}(\cdot)\|_{C^{0}([b_{j}-2\delta,b_{j}])}.$
@@
$\|d_{j,\pm 1}(\cdot;a_{j})\|_{C^{1}([a_{j},b_{j}])}$ $\leq C\|f_{j}(\cdot;a_{j})-f_{j}(\cdot)\|_{C^{0,1}([a_{j},a_{j}+2\delta])},$
$\|d_{j,\pm 1}(\cdot;b_{j})\|_{C^{1}([a_{j},b_{j}])}$ $\leq C\|f_{j}(\cdot;b_{j})-f_{j}(\cdot)\|_{C^{0,1}([b_{j}-2\delta,b_{j}])}.$
@@
$\mathbf{T}_{n}(x,y)=\mathbf{S}_{n}(x,y)\begin{cases}\begin{bmatrix}d_{j,1}^{\sharp}(x,y;c)+1&0\\ 0&d_{j,-1}^{\sharp}(x,y;c)+1\end{bmatrix}&z\in\Omega_{j,\pm},~{}~{}|z-c|<\epsilon,~{}~{}c=a_{j},b_{j},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
We find that $\mathbf{T}_{n}(x,y)$ solves the hybrid $\overline{\partial}$ problem for $c=a_{j},b_{j}$
$\mathbf{T}_{n}^{+}(x,y)$ $=\mathbf{T}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&1\end{bmatrix}&z\in\Sigma_{j},\vspace{3pt}\\ \begin{bmatrix}0&w_{j}(x)/f_{j}(x;c)\\ -f_{j}(x;c)/w_{j}(x)&0\end{bmatrix}&|z-c|<\delta,~{}~{}z\in(a_{j},b_{j}),\vspace{3pt}\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&|z-c|>2\delta,~{}~{}z\in(a_{j},b_{j}),\vspace{3pt}\\ \begin{bmatrix}0&\rho(x)\frac{d_{j,-1}^{\sharp}(x,y;c)+1}{d_{j,1}^{\sharp}(x,y;c)+1}\\ -\frac{d_{j,1}^{\sharp}(x,y;c)+1}{d_{j,-1}^{\sharp}(x,y;c)+1}\rho(x)^{-1}&0\end{bmatrix}&\delta\leq|z-c|\leq 2\delta,~{}~{}z\in(a_{j},b_{j}),\vspace{3pt}\\ \,\mathrm{e}^{-\Delta_{j}{\boldsymbol{\sigma}_{3}}}&z\in(b_{j},a_{j+1}),\end{cases}$
¯$\overline{\partial}\mathbf{T}_{n}(x,y)$ $=\mathbf{T}_{n}(x,y)\mathbf{W}_{n}(x,y),$
$\mathbf{W}_{n}(x,y)$ ={[¯$=\begin{cases}\begin{bmatrix}\frac{\overline{\partial}d_{j,1}(x,y;c)}{d_{j,1}^{\sharp}(x,y;c)+1}&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}\frac{d_{j,1}^{\sharp}(x,y;c)+1}{d_{j,-1}^{\sharp}(x,y;c)+1}&\frac{\overline{\partial}d_{j,-1}^{\sharp}(x,y;c)}{d_{j,-1}^{\sharp}(x,y;c)+1}\end{bmatrix}&z\in\Omega_{j,+},~{}~{}|z-c|<\epsilon,\vspace{3pt}\\ \vspace{3pt}\begin{bmatrix}\frac{\overline{\partial}d_{j,1}^{\sharp}(x,y;c)}{d_{j,1}^{\sharp}(x,y;c)+1}&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}\frac{d_{j,1}^{\sharp}(x,y;c)+1}{d_{j,-1}^{\sharp}(x,y;c)+1}&\frac{\overline{\partial}d_{j,-1}(x,y;c)}{d_{j,-1}^{\sharp}(x,y;c)+1}\end{bmatrix}&z\in\Omega_{j,-}~{}~{}|z-c|<\epsilon,\\ \begin{bmatrix}0&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&0\end{bmatrix}&z\in\Omega_{j,+},~{}~{}|z-a_{j}|>\epsilon,~{}~{}|z-b_{j}|>\epsilon,\vspace{3pt}\\ \vspace{3pt}\begin{bmatrix}0&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&0\end{bmatrix}&z\in\Omega_{j,-},~{}~{}|z-a_{j}|>\epsilon,~{}~{}|z-b_{j}|>\epsilon,\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}),\end{cases}$
@@
$\tilde{\rho}(x;\delta):=\begin{cases}\rho(x;c)&|x-c|<\delta,~{}~{}c=a_{j},b_{j},\\ \rho(x)&|x-c|>2\delta,~{}~{}c=a_{1},b_{1},\ldots,a_{g+1},b_{g+1},\\ \rho(x)\frac{d_{j,-1}^{\sharp}(x,y;c)+1}{d_{j,1}^{\sharp}(x,y;c)+1}&\delta\leq|x-c|\leq 2\delta,~{}~{}c=a_{j},b_{j}.\end{cases}$

### v4

@@ 2.1. The first $\overline{\partial}$ deformation of the FIK RH problem
We will use this factorization locally near each interval $a_{j},b_{j}]$. Define
$\mathbf{Z}_{n}(x,y)=\mathbf{Y}_{n}(z;\mu)\begin{cases}\begin{bmatrix}1&0\\ -f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,+},\vskip 3.0pt\\ \vskip 3.0pt\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,-},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
@@
$\overline{\partial}\mathbf{Z}_{n}(x,y)=\begin{cases}\mathbf{Y}_{n}(z;\mu)\begin{bmatrix}0&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,+},\vskip 3.0pt\\ \vskip 3.0pt\mathbf{Y}_{n}(z;\mu)\begin{bmatrix}0&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,-},\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}).\end{cases}$
@@
$\overline{\partial}\mathbf{Z}_{n}(x,y)=\begin{cases}\mathbf{Z}_{n}(z;\mu)\begin{bmatrix}0&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,+},\vskip 3.0pt\\ \vskip 3.0pt\mathbf{Z}_{n}(z;\mu)\begin{bmatrix}0&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,-},\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}).\end{cases}$
@@
$\mathbf{Z}_{n}^{+}(x,y)=\mathbf{Z}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j},\vskip 3.0pt\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in[a_{j},b_{j}].\end{cases}$
@@
We find that $\mathbf{S}_{n}(x,y)$ solves the normalized hybrid RH-$\overline{\partial}$ problem
$\mathbf{S}_{n}^{+}(x,y)$ $=\mathbf{S}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ \,\mathrm{e}^{-2n\mathfrak{g}(z)}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j},\vskip 3.0pt\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in(a_{j},b_{j}),\vskip 3.0pt\\ \,\mathrm{e}^{-n\Delta_{j}{\boldsymbol{\sigma}_{3}}}&z\in(b_{j},a_{j+1}),\end{cases}$
$\overline{\partial}\mathbf{S}_{n}(x,y)$ $=\mathbf{S}_{n}(z)\begin{cases}\begin{bmatrix}0&0\\ -\,\mathrm{e}^{-2n\mathfrak{g}(z)}\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,+},\vskip 3.0pt\\ \vskip 3.0pt\begin{bmatrix}0&0\\ \,\mathrm{e}^{-2n\mathfrak{g}(z)}\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,-},\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}),\end{cases}$
@@
We recall here that $\mathfrak{g}^{+}(x)+\mathfrak{g}^{-}(x)=0$ for $x\in[a_{j},b_{j}]$ and $\mathfrak{g}^{+}(x)-\mathfrak{g}_{-}(x)={\Delta}_{j}$ for $x\in[b_{j},a_{j+1}]$ where $\Delta_{j}$ is a constant.
@@ 2.2. The second $\overline{\partial}$ deformation of the FIK RH problem
$d_{j,-1}(x;c)$ $=\left[\sqrt{\frac{f_{j}(x;c)}{f_{j}(x)}}-1\right]{(1-b_{\tau}(x;c))},\quad c=a_{j},b_{j},$
$d_{j,+}(x;c)$ $=\left[\sqrt{\frac{f_{j}(x)}{f_{j}(x;c)}}-1\right]{(1-b_{\tau}(x;c))},\quad c=a_{j},b_{j},$
and set $d_{j,\pm}^{\sharp}(x,y;c)$ to be the extension, as in Appendix B, with $\tau=\delta<\epsilon/4$ and $\kappa=0$$d_{j,\pm}^{\sharp}(x,y;c)=d_{j,\pm}(x;c)$ for $|y|<\epsilon/4$. More explicitly, for example,
$\begin{split}d_{j,\pm}^{\sharp}(x,y;a_{j})&=(1-b_{\delta}(|y|;0))\left[\tilde{b}(x)d_{j,\pm}(x;a_{j})\right.\\ &\left.+(1-b_{\delta}(x;a_{j})(1-\beta(x-a_{j},y))d_{j,\pm}(x;a_{j})+\beta(x-a,y)(1-b_{\delta}(x;a_{j}))d_{j,\pm}(a_{j};a_{j})\right].\end{split}$
We write
@@
for $\xi$ between $f_{j}(x;c)$ and $f_{j}(x)$. Since both of these functions are bounded below on $0,2\delta]$ we find
$\|d_{j,\pm}(\cdot;a_{j})\|_{C^{0}([a_{j},b_{j}])}$ $\leq C\|f_{j}(\cdot;a_{j})-f_{j}(\cdot)\|_{C^{0}([a_{j},a_{j}+2\delta])},$
$\|d_{j,\pm}(\cdot;b_{j})\|_{C^{0}([a_{j},b_{j}])}$ $\leq C\|f_{j}(\cdot;a_{j})-f_{j}(\cdot)\|_{C^{0}([b_{j}-2\delta,b_{j}])}.$
@@
$\|d_{j,\pm}(\cdot;a_{j})\|_{C^{1}([a_{j},b_{j}])}$ $\leq C\|f_{j}(\cdot;a_{j})-f_{j}(\cdot)\|_{C^{0,1}([a_{j},a_{j}+2\delta])},$
$\|d_{j,\pm}(\cdot;b_{j})\|_{C^{1}([a_{j},b_{j}])}$ $\leq C\|f_{j}(\cdot;b_{j})-f_{j}(\cdot)\|_{C^{0,1}([b_{j}-2\delta,b_{j}])}.$
@@
$\mathbf{T}_{n}(x,y)=\mathbf{S}_{n}(x,y)\begin{cases}\begin{bmatrix}d_{j,{-}}^{\sharp}(x,y;c)+1&0\\ 0&d_{j,{+}}^{\sharp}(x,y;c)+1\end{bmatrix}&z\in\Omega_{j,\pm},~~|z-c|<\epsilon,~~c=a_{j},b_{j},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
We find that $\mathbf{T}_{n}(x,y)$ solves the hybrid RH-$\overline{\partial}$ problem for $c=a_{j},b_{j}$
$\mathbf{T}_{n}^{+}(x,y)$ $=\mathbf{T}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ \,\mathrm{e}^{-2n\mathfrak{g}(z)}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j},\vskip 3.0pt\\ \begin{bmatrix}0&w_{j}(x)/f_{j}(x;c)\\ -f_{j}(x;c)/w_{j}(x)&0\end{bmatrix}&|z-c|<\delta,~~z\in(a_{j},b_{j}),\vskip 3.0pt\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&|z-c|>2\delta,~~z\in(a_{j},b_{j}),\vskip 3.0pt\\ \begin{bmatrix}0&\rho(x)\frac{d_{j,-}^{\sharp}(x,y;c)+1}{d_{j,+}^{\sharp}(x,y;c)+1}\\ -\frac{d_{j,+}^{\sharp}(x,y;c)+1}{d_{j,-}^{\sharp}(x,y;c)+1}\rho(x)^{-1}&0\end{bmatrix}&\delta\leq|z-c|\leq 2\delta,~~z\in(a_{j},b_{j}),\vskip 3.0pt\\ \,\mathrm{e}^{-\Delta_{j}{\boldsymbol{\sigma}_{3}}}&z\in(b_{j},a_{j+1}),\end{cases}$
$\overline{\partial}\mathbf{T}_{n}(x,y)$ $=\mathbf{T}_{n}(x,y)\mathbf{W}_{n}(x,y),$
$\mathbf{W}_{n}(x,y)$ $=\begin{cases}\begin{bmatrix}\frac{\overline{\partial}d_{j,+}^{\sharp}(x,y;c)}{d_{j,+}^{\sharp}(x,y;c)+1}&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}\frac{d_{j,+}^{\sharp}(x,y;c)+1}{d_{j,-}^{\sharp}(x,y;c)+1}&\frac{\overline{\partial}d_{j,-}^{\sharp}(x,y;c)}{d_{j,-}^{\sharp}(x,y;c)+1}\end{bmatrix}&z\in\Omega_{j,+},~~|z-c|<\epsilon,\vskip 3.0pt\\ \vskip 3.0pt\begin{bmatrix}\frac{\overline{\partial}d_{j,+}^{\sharp}(x,y;c)}{d_{j,+}^{\sharp}(x,y;c)+1}&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}\frac{d_{j,+}^{\sharp}(x,y;c)+1}{d_{j,-}^{\sharp}(x,y;c)+1}&\frac{\overline{\partial}d_{j,-}^{\sharp}(x,y;c)}{d_{j,-}^{\sharp}(x,y;c)+1}\end{bmatrix}&z\in\Omega_{j,-}~~|z-c|<\epsilon,\\ \begin{bmatrix}0&0\\ -\overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&0\end{bmatrix}&z\in\Omega_{j,+},~~|z-a_{j}|>\epsilon,~~|z-b_{j}|>\epsilon,\vskip 3.0pt\\ \vskip 3.0pt\begin{bmatrix}0&0\\ \overline{\partial}f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&0\end{bmatrix}&z\in\Omega_{j,-},~~|z-a_{j}|>\epsilon,~~|z-b_{j}|>\epsilon,\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}),\end{cases}$
@@
$\tilde{\rho}(x;\delta):=\begin{cases}\rho(x;c)&|x-c|<\delta,~~c=a_{j},b_{j},\\ \rho(x)&|x-c|>2\delta,~~c=a_{1},b_{1},\ldots,a_{g+1},b_{g+1},\\ \rho(x)\frac{d_{j,-}^{\sharp}(x,y;c)+1}{d_{j,+}^{\sharp}(x,y;c)+1}&\delta\leq|x-c|\leq 2\delta,~~c=a_{j},b_{j}.\end{cases}$